Tamara has 2 sisters. One of the sisters is 7 years older the Tamar. The other sister is 3 years younger then Tamara. The product of her sisters ages is 24. How old is Tamra
Answer by ankor@dixie-net.com(22740)

Tamara has 2 sisters. One of the sisters is 7 years older the Tamar.
The other sister is 3 years younger then Tamara.
The product of her sisters ages is 24. How old is Tamra
:
Let t = T's age
then
(t+7) = older sister's age
and
(t-3) = younger sister's age
:
The product of her sisters ages is 24.
(t+7)(t-3) = 24
FOIL
t^2 - 3t + 7t - 21 = 24
:
t^2 + 4t - 21 - 24 = 0
:
t^2 + 4t - 45 = 0
factors to
(t+9)(t-5) = 0
Positive solution
t = 5 yrs is T's age
;
:
Check solution by find the product of the sisters age
(5+7) * (5-3) = 24